OpenStudy (anonymous):

thank you. i have another question: a jar contains 1000 identical nails. If a number of nails weighing a total of 1500g were removed from the jar, there would be 850 nails left in the jar. How many g does 1 nail wiegh?

OpenStudy (phi):

Can you make up some variables?

OpenStudy (anonymous):

1000 nails-1500g= 850 nails

OpenStudy (phi):

you are mixing apple and oranges or in this case nails and grams?

OpenStudy (anonymous):

i guess. im very confused.

OpenStudy (phi):

I would take each sentence and try to simplify it. a jar contains 1000 identical nails. I would say T= 1000 T for total If a number of nails weighing a total of 1500g a number of nails -- This sounds like an unknown. call it X weighs 1500 g we know each nail will weigh something. What do we call it? Y (why not!) X*Y= 1500

OpenStudy (phi):

there would be 850 nails we started with T, took away X and are left with 850 T-X= 850, or using 1000 for T (simplify) 1000-X= 850

OpenStudy (phi):

How many g does 1 nail weigh? we (I) called this Y. so we have to find Y. can you summarize the equations we have?

OpenStudy (anonymous):

1000-x=850 x=150

OpenStudy (phi):

and don't forget X*Y= 1500 g

OpenStudy (anonymous):

y=10g

OpenStudy (phi):

Do you see how I did it? Write stuff down, and see if you can make equations. Try to make sense of the mess....